# PACE: Pose Annotations in Cluttered Environments

### v1

@@ Abstract
Pose estimation is a crucial task in computer vision, enabling tracking and manipulating objects in images or videos. While several datasets exist for pose estimation, there is a lack of large-scale datasets specifically focusing on cluttered scenes with occlusions. This limitation is a bottleneck in the development and evaluation of pose estimation methods, particularly toward the goal of real-world application in environments where occlusions are common. Addressing this, we introduce PACE (Pose Annotations in Cluttered Environments), a large-scale benchmark designed to advance the development and evaluation of pose estimation methods in cluttered scenarios. PACE encompasses 54,945 frames with 257,673 annotations across 300 videos, covering 576 objects from 44 categories and featuring a mix of rigid and articulated items in cluttered scenes. To annotate the real-world data efficiently, we developed an innovative annotation system utilizing a calibrated 3-camera setup. We test state-of-the-art algorithms in PACE along two tracks: pose estimation, and object pose tracking, revealing the benchmark’s challenges and research opportunities. We plan to release PACE as a public evaluation benchmark, along the annotations tools we developed, to stimulate further advancements in the field. Our code and data is available on https://github.com/qq456cvb/PACE.
@@ 1 Introduction
The field of 3D object pose estimation is integral to a myriad of applications, particularly within robotic manipulation. Recent advancements in both instance and category-level pose estimation have been significant, bolstered by deep learning approaches, and perhaps more importantly, data.
@@
PoseCNN [45] advanced pose estimation into the deep learning era, and simultaneously introduced the influential YCB-Video dataset. This benchmark has catalyzed methodological development and offered a consistent evaluation platform. Additionally, the Benchmark for 6D Object Pose Estimation (BOP) challenges have consolidated datasets and refined evaluation metrics for instance-level pose estimation.
@@
In parallel, the NOCS [41] dataset has addressed category-level pose estimation, albeit with a smaller real-world dataset for validation and testing. Despite these strides, the field grapples with a fundamental challenge: existing evaluation datasets are too constrained to thoroughly benchmark the capabilities of pose estimation algorithms. The NOCS REAL275 dataset, for instance, spans only six categories and includes a mere 18 videos. This limitation has led to a performance saturation on current benchmarks, as depicted in Table 1, rendering it ambiguous whether methods are getting tuned to the dataset or improving in general.
| Bottle | Bowl | Camera | Can | Laptop | Mug | |
| SAR-Net [24] | 54.0 | 66.0 | 0.4 | 62.2 | 50.0 | 21.2 |
| HS-Pose [47] | 51.1 | 94.7 | 3.9 | 75.4 | 85.2 | 26.4 |
@@
In this work, we introduce PACE (Common Objects with Pose Annotations), a benchmark for pose estimation, and present a comprehensive study evaluating a wide range of pose estimation and tracking methods. Our contributions are threefold:
@@
The PACE dataset: This dataset includes 576 objects across 44 categories, captured in 300 video clips within diverse scenes. With an average of 183 frames per clip, the dataset encompasses 54,945 frames and 257,673 annotations, providing a large-scale benchmark for pose estimation.
@@
Our evaluation study: To the best of our knowledge, we are the first to analyze and report the performance of state-of-the-art pose estimation methods in a large-scale cluttered setting. These results provide valuable insights that the scalability and generalizability of state-of-the-art methods, making it clear that they are far from reliable in general.
@@
Our annotation pipeline: We will open-source our annotation pipeline, which harnesses a calibrated 3-camera system, enhancing the precision and scalability of annotating poses in real data. This tool significantly mitigates human error and reduces the effort required for annotating 3D poses, providing a solution to one of the major bottlenecks in pose dataset creation.
@@
Overall, our work aims to propel the development of more robust and generalizable pose estimation techniques, thereby facilitating progress towards successful pose estimation in the real world.
| Modality | Cat. | Obj. | Vid. | Img. | Anno. | CAD | Dyn. | Occ. | Marker-free | Artic. | Piled | |
| YCB-Video [45] | RGBD | 1 | 21 | 12 | 20K | 99K | ✓ | ✗ | ✓ | ✓ | ✗ | ✓ |
| LINEMOD-O [3] | RGBD | 1 | 8 | 1 | 1.2K | 9.2K | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ |
| NAVI [16] | RGBD | 1 | 36 | 324 | 10K | 10k | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
@@
| PACE (Ours) | RGBD | 44 | 576 | 300 | 55K | 258K | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
2 Related Works
2.1 Object Pose Datasets
Instance-Level Pose Datasets
YCB-Video dataset [45] is a comprehensive resource for 6D object pose estimation, containing a large number of video frames with accurate pose annotations for 21 objects. LINEMOD-Occluded dataset [3] offers a challenging setting for pose estimation with piled multiple objects in occluded scenes. NAVI dataset [16] presents casually captured images of objects with high-quality 3D scans and precise 2D-3D alignments for advanced 3D reconstruction tasks.
Category-Level Pose Datasets
CO3D [32] offers 1.5 million frames from nearly 19,000 videos across 50 MS-COCO categories for category-specific 3D reconstruction and view synthesis. The Scan2CAD [2] dataset aligns 14225 CAD models from ShapeNet to 1506 ScanNet scans, promoting CAD model alignment in RGB-D scans. Pix3D [36] is a benchmark with image-shape pairs and pixel-level 2D-3D alignment, aiding in shape reconstruction and retrieval. The HOI-4D dataset [29], with 2.4M RGB-D frames over 4000 sequences, enables research in category-level human-object interaction. HANDAL [11] focuses on pose estimation and affordance prediction for robotics-ready manipulable objects. A comparative comparison with other datasets is in Table 2.
@@ 2.2 Pose Estimation Methods
Instance-level Pose Estimation
PPF (Point Pair Features) [8] set the pre-deep learning standard for instance-level pose estimation, utilizing local geometric features from point clouds. The deep learning era began with PoseCNN [45], leading to several advanced methods. DeepIM [23] approached pose estimation as an image matching task, iteratively refining estimations. DenseFusion [38] combined global and local features for pose estimation in cluttered scenes, while CosyPose [19] integrated a global refinement strategy in its end-to-end pipeline. SurfEmb [12] leveraged surface embeddings for correspondence matching, and GDRNPP [40] employed geometry-guided regression for enhanced prediction.
Category-level Pose Estimation
Category-level pose estimation extends the challenge to generic object categories. NOCS [41] introduced a unified coordinate space for all objects, predicting object NOCS maps from RGB images. SGPA [4] aims to adapt the structure-guided prior in the pose estimation process, while SAR-Net [24] using shape alignment and symmetric correspondence to estimate a coarse 3D object shape and facilitate object center and size estimation. Recently, HS-Pose [47] proposes a network structure with an HS-layer that extends 3D graph convolution to extract hybrid scope latent features from point clouds for category-level object pose estimation.
@@ 2.3 Pose Tracking Methods
Instance-level Pose Tracking Methods like RBOT [37] use RGB data and 3D models to track multiple objects, employing color histograms in their cost function. PoseRBPF [5] separates rotation and translation, using an autoencoder for rotation feature embeddings. ICG [34] iteratively refines pose using geometric cues and is effective for textureless objects, with extensions incorporating visual data [35]. The first deep learning tracker, D6DT [9], and se(3)-TrackNet [43] predict frame-to-frame relative poses, using a render-and-compare strategy.
Category-level Pose Tracking
6-PACK [39] marks the onset of category-level tracking, using DenseFusion [38] features and an attention mechanism for unsupervised keypoint ordering and interframe motion via keypoint matching. CenterPoseTrack [26] projects 2D keypoints from 3D bounding box vertices, achieving RGB-based scale-invariant tracking. BundleTrack [42] generalizes pose tracking without relying on 3D models, instead using video segmentation and pose graph optimization. CAPTRA [44] tracks 9DoF poses for rigid and articulated objects, with subnetworks for rotation regression and normalized coordinate prediction, facilitating analytical 3D size and translation computation.
@@ 3 Construction of PACE
3.1 Acquisition of 3D Common Object Scans
We begin by digitizing an extensive collection of commonplace objects. These items are categorized into 44 distinct classes, as represented in Figure 3.
The Einscan Pro 2X is utilized for rapid scanning of all objects, typically completing within 5 to 10 minutes per object. To expedite this process, we employ a rotatable platform to acquire multiple viewpoints of the objects. After scanning, objects are manually aligned to a standard pose within a uniform coordinate system. We center its axis-aligned bounding box at the origin and align the bounding box orientation along a common axis within the category. We annotate rotational symmetries with the corresponding rotation matrices. High-resolution meshes are then simplified to lower-resolution for smoother annotation workflows.
@@ 3.1 Acquisition of 3D Common Object Scans
Diverging from many prior pose estimation datasets, our collection encompasses a very wide set of objects, including articulated objects from the AKB48 [27] dataset, namely: scissors, cutters, clips, and boxes. We adopt the alignment methodology from the original AKB48 dataset, without modification. These objects are segmented into multiple parts with hierarchical relationships, presenting a nuanced challenge for pose estimation.
@@ 3.2 RGB-D Sequence Acquisition
We designed and implemented a 3-camera system to aid in data acquisition and annotation, comprising three Intel Realsense 415 RGB-D cameras affixed to a metal framework, as illustrated in the bottom of Figure 2. The advantages of this setup include:
Tripling the data yield.
Reducing ambiguity in pose annotation, especially regarding translation along the depth dimension, by using multi-view imagery to enforce consistency across all views.
Enhancing tracking accuracy of static objects with Aruco markers from all three views, making PnP more stable.
@@
We calibrate this multi-camera system through a semi-automatic process. Aruco markers initially intended for calibration proved insufficient for high-accuracy rotation estimation. Hence, we resorted to trifocal tensor estimation, i.e. TFT [17]. The process begins with feature extraction and matching, followed by bundle adjustment to refine the positions of the 3D landmarks and camera poses. For reliable feature matching, we employ the SuperPoint [7] descriptor and SuperGlue [33] matcher, using a stringent threshold for matching. We observe that rotational component of the resulting extrinsic parameters is generally precise, but the translation aspect suffers from scale ambiguity inherent in Structure-from-Motion approaches. We correct this by calibrating the scale against markers, applying the following formula to obtain the optimal scale factor:
where is the TFT predicted translation (up to a scale) from camera to camera , is the marker-calibrated translation in real metric scale. We set the extrinsics of the first camera to be the identity matrix.
In cases of repetitive texture patterns, manual intervention is required to establish reliable feature correspondences due to the limitations of SuperPoint+SuperGlue.
@@ 3.3 Annotation of Pose Ground-Truths
Previous methodologies have utilized Aruco markers to automate pose estimation through Perspective-n-Points; however, this approach has two drawbacks:
@@
Marker-based annotations are inapplicable to dynamic objects, thus restricting the method’s utility to static scenarios.
Annotation of Static Object Poses
To address the first issue, we employ markers to automate the annotation of static object poses, and then remove the marker appearances from the dataset. We achieve this with a marker inpainting strategy, detailed as follows. Initially at step 1, we place a marker (Marker 1) somewhere within the camera’s field of view. We then record a short video with this marker in view. In step 2, we place a second marker (Marker 2) at a chosen distance from the first, and record another video. After step 2, we remove Marker 1, and begin the actual object capture process (with only Marker 2 present). After this process, we end up with: (1) frames with Marker 1 only, which clearly depict the surface where Marker 2 will later appear; (2) frames capturing both markers, providing helpful calibration cues; (3) frames with Marker 2 only, which represent our main capture. We use the frames from the first two steps to seamlessly inpaint [30] Marker 2’s area in the main capture, as depicted in Figure 4. We leverage Marker 2 for automated pose tracking, and manually correct the tracking every 40 frames in case of drift.
Annotation of Dynamic Object Poses
For dynamic objects, the traditional marker-based tracking approach is insufficient, as the scene markers do not move in tandem with the objects. In such instances, we harness the capabilities of BundleTrack [42], an advanced RGB image-based tracking algorithm. BundleTrack conducts feature correspondence analysis between successive frames to estimate poses, complemented by a bundle adjustment algorithm to optimize keyframes globally and minimize tracking errors. Despite BundleTrack’s proficiency in approximating poses, it is prone to drift, necessitating the manual adjustment of poses every ten frames to ensure precision. This delicate step represents the most labor-intensive aspect of our annotation pipeline, given the current limitations of state-of-the-art tracking techniques in complex scenarios.
@@
Upon successful pose annotation, we generate occlusion-aware segmentation masks, using z-buffer rendering techniques. In cases where hand interactions are involved, we employ the SAM [18] model to delineate hand masks, and subsequently subtract these from the previously computed object masks to achieve accurate segmentation.

### v3

# PACE: A Large-Scale Dataset with Pose Annotations in Cluttered Environments

@@ Abstract
We introduce PACE (Pose Annotations in Cluttered Environments), a large-scale benchmark designed to advance the development and evaluation of pose estimation methods in cluttered scenarios. PACE provides a large-scale real-world benchmark for both instance-level and category-level settings. The benchmark consists of 55K frames with 258K annotations across 300 videos, covering 238 objects from 43 categories and featuring a mix of rigid and articulated items in cluttered scenes. To annotate the real-world data efficiently, we develop an innovative annotation system with a calibrated 3-camera setup. Additionally, we offer PACE-Sim, which contains 100K photo-realistic simulated frames with 2.4M annotations across 931 objects. We test state-of-the-art algorithms in PACE along two tracks: pose estimation, and object pose tracking, revealing the benchmark’s challenges and research opportunities. Our benchmark code and data is available on https://github.com/qq456cvb/PACE.
@@ 1 Introduction
The field of 3D object pose estimation is integral to a myriad of applications, particularly within robotic manipulation. Recent advancements in both instance and category-level pose estimation have been bolstered by deep learning approaches, and perhaps more importantly, by data.
@@
PoseCNN [38] advanced pose estimation into the deep learning era, and simultaneously introduced the influential YCB-Video dataset. This benchmark has catalyzed methodological development and offered a consistent evaluation platform. Besides, the Benchmark for 6D Object Pose Estimation (BOP) challenges [14] have consolidated datasets metrics for instance-level pose estimation.
@@
In parallel, NOCS [34] has addressed category-level pose estimation, albeit with a smaller real-world dataset for validation and testing. Despite these strides, the field grapples with a fundamental challenge: state-of-the-art pose estimation models still perform poorly in real-world settings, and existing evaluation datasets are too constrained to thoroughly reveal this fact. The NOCS-REAL275 dataset, for instance, spans only six categories and includes a mere 18 videos. This limitation has led to an array of published models which perform near-perfectly on NOCS but do not generalize to new data, as demonstrated in Figure 2.
@@
In this work, we introduce PACE, a benchmark for pose estimation, and present a comprehensive study evaluating a wide range of pose estimation and tracking methods. Our contributions are threefold:
@@
The PACE dataset: This dataset includes 238 objects across 43 categories, captured in 300 video clips within diverse scenes. With an average of 183 frames per clip, the dataset consists of 55K frames and 258K annotations, providing a large-scale benchmark for pose estimation. We also provide PACE-Sim, a photo-realistic training dataset which contains 100K frames with 2.4M annotations across 931 objects.
@@
Our evaluation study: To the best of our knowledge, we are the first to analyze and report the performance of state-of-the-art (SOTA) pose estimation methods in a large-scale cluttered setting. These results provide valuable insights on the scalability and generalizability of SOTA methods, making it clear that they are far from reliable in general.
@@
Our annotation pipeline: We will open-source our annotation pipeline, which uses a calibrated 3-camera system, enhancing the precision and scalability of annotating poses in real data. This tool significantly mitigates human error and reduces the effort required for annotating 3D poses, providing a solution to one of the major bottlenecks in pose dataset creation.
@@
Overall, our work aims to support the development of more robust and generalizable pose estimation techniques, thereby facilitating progress towards successful pose estimation in the real world.
2 Related Works
2.1 Object Pose Datasets
2.1.1 Instance-Level Pose Datasets
YCB-Video dataset [38] is a comprehensive resource for 6D object pose estimation, containing a large number of video frames with accurate pose annotations for 21 objects. LINEMOD-Occluded [3] offers a challenging setting for pose estimation with piled multiple objects in occluded scenes. NAVI dataset [15] presents casually captured images of objects with high-quality 3D scans and precise 2D-3D alignments for 3D reconstruction tasks. Recently, StereoObj-1M [24] is proposed to address challenging cases such as object transparency, in addition to common challenges of occlusion and symmetry.
2.1.2 Category-Level Pose Datasets
NOCS-REAL275 [34] and Wild6D [40] both annotate a limited number of categories with upright poses. Objectron [1] is a collection of short, single-object-centric video clips with no clutters. The Scan2CAD [2] dataset aligns 14225 CAD models from ShapeNet to 1506 ScanNet scans, promoting CAD model alignment in RGB-D scans. Pix3D [29] is a benchmark with image-shape pairs and pixel-level 2D-3D alignment, aiding in shape reconstruction and retrieval. HANDAL [11] focuses on pose estimation and affordance prediction for robotics-ready manipulable objects. HouseCat6D [17] encompasses 194 diverse objects across 10 household categories with additional grasp annotations. ROPE [41] provides a large-scale evaluation benchmark covering a wide spectrum of categories. However, it does not include challenging scenarios such as moving and piled objects. Table 1 presents a comparative comparison of these datasets, alongside our new dataset PACE.
| Modality | Cat. | Obj. | Vid. | Img. | Anno. | CAD | Mov. | Occ. | Marker-free | Artic. | Piled | |
| YCB-Video [38] | RGBD | - | 21 | 12 | 20K | 99K | ✓ | ✗ | ✓ | ✓ | ✗ | ✓ |
| LINEMOD-O [3] | RGBD | - | 8 | 1 | 1.2K | 9.2K | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ |
| NAVI [15] | RGBD | - | 36 | 324 | 10K | 10k | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| StereoObj-1M [24] | RGBD | - | 18 | 182 | 393K | 1.5M | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ |
@@ 2.1.2 Category-Level Pose Datasets
| HouseCat6D [17] | RGBD | 10 | 192 | 41 | 24K | 160K | ✓ | ✗ | ✓ | ✓ | ✗ | ✓ |
| ROPE [41] | RGBD | 149 | 581 | 363 | 332K | 1.5M | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| PACE (Ours) | RGBD | 43 | 238 | 300 | 55K | 258K | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@ 2.2 Pose Estimation Methods
2.2.1 Instance-level Pose Estimation
PPF (Point Pair Features) [8] is a pre-deep learning standard for instance-level pose estimation, using local geometric features from point clouds. PoseCNN [38] is the first model trained end-to-end for the task. CosyPose [19] integrates a global refinement strategy in its end-to-end pipeline. SurfEmb [12] leverages surface embeddings for correspondence matching, and GDRNPP [33] uses geometry-guided regression.
2.2.2 Category-level Pose Estimation
Category-level pose estimation extends the challenge to generic object categories. NOCS [34] introduces a unified coordinate space for all objects, predicting object NOCS maps from RGB images. SGPA [4] aims to adapt the structure-guided prior in the pose estimation process, while SAR-Net [21] uses shape alignment and symmetric correspondence to estimate a coarse 3D object shape and facilitate object center and size estimation. Recently, HS-Pose [43] proposes a network structure with a HS-layer that extends 3D graph convolution to extract hybrid scope latent features from point clouds for category-level object pose estimation. GenPose [42] incorporates probabilistic stable diffusion models into pose estimation, achieving strong performance.
@@ 2.3 Pose Tracking Methods
2.3.1 Instance-level Pose Tracking
Methods like RBOT [30] use RGB data and 3D models to track multiple objects, employing color histograms in their cost function. PoseRBPF [5] separates rotation and translation, using an autoencoder for rotation feature embeddings. ICG [28] iteratively refines pose using geometric cues and is effective for textureless objects, with extensions incorporating visual data [27]. The first deep learning tracker, D6DT [9], and se(3)-TrackNet [36] predict frame-to-frame relative poses, using a render-and-compare strategy.
2.3.2 Category-level Pose Tracking
6-PACK [31] performs category-level tracking, using DenseFusion [32] features and an attention mechanism for unsupervised keypoint ordering and interframe motion via keypoint matching. BundleTrack [35] performs pose tracking without relying on 3D models, instead using video segmentation and pose graph optimization. CAPTRA [37] tracks 9DoF poses for rigid and articulated objects, with heads for rotation regression and normalized coordinate prediction, from which 3D size and translation can be computed.
@@ 3 Construction of PACE
3.1 Acquisition of 3D Scans of Common Objects
We began by digitizing an extensive collection of commonplace objects. These items were categorized into 43 distinct categories, as represented in Figure 5. The Einscan Pro 2X was utilized for rapid scanning of all objects, typically completing within 5 to 10 minutes per object. To expedite this process, we employed a rotatable platform to acquire multiple viewpoints of objects. After scanning, objects were manually aligned to a standard pose within a uniform coordinate system. For each object, we centered its axis-aligned bounding box at the origin and then aligned the bounding box orientation along a common axis within the category. We additionally annotated rotational symmetries. High-resolution meshes were simplified to lower-resolution ones for a smoother annotation workflow.
@@ 3.1 Acquisition of 3D Scans of Common Objects
Diverging from many prior pose estimation datasets, our collection encompasses a very wide set of objects, including articulated objects from the AKB48 [23] dataset, namely: scissors, cutters, clips, and boxes. We adopt the alignment methodology from the original AKB48 dataset, without modification. These objects are segmented into multiple parts with hierarchical relationships, presenting a difficult challenge for pose estimation.
@@ 3.2 RGB-D Sequence Acquisition
We implemented a 3-camera system to aid in data acquisition and annotation, comprising three Intel Realsense D415 RGB-D cameras affixed to a metal framework, as illustrated in the bottom of Figure 3. The advantages of this setup include: (1) tripling the data yield; (2) reducing ambiguity in pose annotation, especially regarding translation along the depth dimension, by enforcing annotation consistency across all 3 views; (3) enhancing tracking accuracy of still objects with ArUco markers [10], making PnP more stable.
@@
We calibrated this multi-camera system through a semi-automatic process. ArUco markers [10] alone proved insufficient for high-accuracy rotation estimation. Hence, we resorted to trifocal tensor estimation, i.e. TFT [16]. The process begins with feature extraction and matching, followed by bundle adjustment to refine the positions of the 3D landmarks and camera poses. For reliable feature matching, we employed the SuperPoint [7] descriptor and SuperGlue [26] matcher, using a stringent threshold for matching. We observed that rotational component of the resulting extrinsic parameters is precise, but the translation aspect suffers from scale ambiguity inherent in Structure-from-Motion approaches. We corrected this by calibrating the scale against markers, applying the following formula to obtain the optimal scale factor, by setting gradient of to zero:
where is the TFT predicted translation (up to a scale) from camera to camera , is the marker-calibrated translation in real metric scale. We set the extrinsics of the first camera to be the identity matrix. We use Intel’s dynamic calibration tool to calibrate intrinsics. In cases of repetitive textures, where SuperPoint+SuperGlue were unreliable, we made correspondences manually.
@@ 3.3 Annotation of Pose Ground-Truths
Previous works have used ArUco markers to automate pose estimation through Perspective-n-Points, and we do this also. Specifically, we used ArUco markers to automate the annotation of still object poses (i.e. not moving), following LINEMOD [3] and NOCS [34]. We note that this step has two limitations:
@@
Marker-based annotations are inapplicable to moving objects (since the markers are typically placed on the background), thus limiting the setup to annotating still objects.
We compensate for both of these issues: we remove marker appearances from the dataset (after using them for annotation), and we introduce additional tools to annotate the moving objects. We describe these compensation steps next.
Marker removal
We removed the marker appearances from the dataset, using a marker inpainting strategy, detailed as follows. In step 1, we place a marker (Marker 1) somewhere within the camera’s field of view. We then record a short video with this marker in view. In step 2, we place a second marker (Marker 2) at a chosen distance from the first, and record another video. In step 3, we remove Marker 1, and begin the actual object capture process (with only Marker 2 present). After this process, we end up with: (1) frames with Marker 1 only, which clearly depict the surface where Marker 2 will later appear; (2) frames capturing both markers, providing helpful calibration cues; (3) frames with Marker 2 only, which represent our main capture. We leverage Marker 2 for automated pose tracking, and manually correct the tracking every 40 frames in case of drift. We use the frames from the first two steps to seamlessly inpaint [25] Marker 2’s area for the final dataset, as depicted in Figure 4.
Annotation of Moving Object Poses
For moving objects, the traditional marker-based tracking approach is insufficient, as the scene markers do not move in tandem with the objects. In such instances, we harnessed the capabilities of BundleTrack [35], an advanced RGB image-based tracking algorithm. BundleTrack conducts feature correspondence analysis between successive frames to estimate poses, complemented by a bundle adjustment algorithm to optimize keyframes globally and minimize tracking errors. Despite BundleTrack’s proficiency in approximating poses, it is prone to drift, necessitating the manual adjustment of poses every ten frames to ensure precision. We adopted a rigorous manual checking process, examining every tenth frame for 3D pointcloud alignment and 2D multi-view agreement. We detail the annotation tool and its graphical interface in the supplementary. This delicate process represents the most labor-intensive aspect of our annotation pipeline.
In order to verify that our annotation process gives accurate poses, we conducted an experiment where we used our setup to manually annotate the poses of Scene 1 from NOCS-REAL275 [34] dataset, and compare the annotations to the provided ground truths. We recorded an average pose annotation error of for rotation and mm for translation.
@@
Upon successful pose annotation, we generated occlusion-aware segmentation masks, via depth rendering. In cases where hand interactions are involved, we employed the segmentation model SAM [18] to estimate hand masks, and subsequently subtracted these from the previously computed object masks to obtain the occlusion-aware segmentation.